OpenStudy (anonymous):

A cylindrical glass is 3 inches in diameter and 6.5 inches tall. How much water can the glass hold when full? Round your answer to two decimal places.

OpenStudy (anonymous):

A. 26.70 cubic inches B. 36.03 cubic inches C. 45.95 cubic inches D. 52.62 cubic inches

OpenStudy (welshfella):

use the formula for the volume of a cylinder V = pi* r^2*h where h = height and r = radius also note - the radius = half of the diameter

OpenStudy (anonymous):

got 183.69

OpenStudy (welshfella):

you have the height h that = 6.5 and rad = 1/2 * 3

OpenStudy (welshfella):

V = pi * (1.5)^2 * 6.5

OpenStudy (anonymous):

i just did the math its 45.95

OpenStudy (welshfella):

yes - correct

OpenStudy (anonymous):

^.^

OpenStudy (anonymous):

Your Welcome XD

OpenStudy (welshfella):

but please don't just give the answer - its against the Code of Conduct

OpenStudy (anonymous):

oh okay no problem.

OpenStudy (welshfella):

mods may ban you for a while...

OpenStudy (anonymous):

>.<